A jeepney follows a circular road with a radius of 300 meters at a speed of 30 m/s. What is the magnitude of the jeepney’s acceleration? A 450 kg jeep taking a turn on a 350 m radius road at 22 m/s. Calculate the centripetal force exerted on the jeep.
1
Expert's answer
2021-10-25T17:54:05-0400
"a=\frac{30^2}{300}=3\frac{m}{s^2}"

"F=450\frac{22^2}{350}=622\ N"